A hospital receives 30% of its COVID-19 vaccine shipments from Ghana Health Service
and the remainder of its shipments from neighbouring West African countries. Each
shipment contains a very large number of vaccine vials. For GHS's shipments, 8% of the
vials are inefective and for the neighbouring countries, 3% of the vials are inefective.
The hospital tests 30 randomly selected vials from a shipment and fi nds that one is
inefective. What is the probability that the shipment came from neighbouring West
African countries?
1
Expert's answer
2020-06-03T17:56:51-0400

The probability that the shipment came from Ghana and neighboring West Africa countries is

"P(G) =0.3, P(G^c) =0.7"

respectively.

The probability of an ineffective vial in Ghana and neighboring countries

"P(I|G) =0.08, P(I|G^c) =0.03"

Given n =30, the probability of getting 1 ineffective vial is

"P(I=1|G)={30\choose 1}0.08^1 0.92^{29}"

=0.2138

"P(I=1|G^c)={30\choose 1}0.03^1 0.97^{29}"

=0.3721

The probability of getting one ineffective vial

"P(I=1) =P(I=1|G) P(G) +P(I=1|G^c) P(G^c)"

=0.2138*0.3+0.3721*0.7

=0.3246

The probability that the shipment came from neighbouring West African countries is calculated by applying the Bayes theorem

"P(G^c|I=1)=\frac {P(I=1|G^c)P(G^c)}{P(I=1)}"

"=\frac {0.3721 *0.7} {0.3246}"

=0.8024